This exercise illustrates just how fast exponential functions grow in the long term. Suppose you start work for a company at age 25. You are offered two rather unlikely retirement options.
Retirement option 1: When you retire, you will receive $15,000 for each year of service.
Retirement option 2: When you start work, the company deposits $2500 into a savings account that pays a monthly rate of 1.3%. When you retire, the account will be closed and the balance given to you.
How much will you have under the second plan at age 55? (Round your answer to the nearest cent.)
$ 600,000

Incorrect: Your answer is incorrect.

How much will you have under the second plan at age 65? (Round your answer to the nearest cent.)
$ 600,000

Incorrect: Your answer is incorrect.
I was wrong on both. Can someone show me how to solve these with the correct answers.


Answer by stanbon(75887)

This exercise illustrates just how fast exponential functions grow in the long term. Suppose you start work for a company at age 25. You are offered two rather unlikely retirement options.
Retirement option 1: When you retire, you will receive $15,000 for each year of service.
Retirement option 2: When you start work, the company deposits $2500 into a savings account that pays a monthly rate of 1.3%. When you retire, the account will be closed and the balance given to you.
How much will you have under the second plan at age 55? (Round your answer to the nearest cent.)
-------
30 years = 20*12 = 240 months
2500*(1.013)^240 = 2500*22.196 = $55489.00
----------------------
How much will you have under the second plan at age 65? (Round your answer to the nearest cent.)
40 years = 40*12 = 480 months
2500*(1.013)^480 = 2500*492.64 = $1,231,611.63
----------
Cheers,
Stan H.
---------------